OpenStudy (anonymous):

Determine the exact values of the intercepts by using the quadratic formula. Then a calculator to evaluate the expressions that you obtain. y=3x^2-12sqrt3x+36

OpenStudy (amistre64):

seems pretty much cut and dry

OpenStudy (amistre64):

which part of the instructions is giving you troubles?

OpenStudy (anonymous):

I am just learning this the quad formula how do i determine a,b,c

OpenStudy (amistre64):

a b and c never change; they are pretty well established as the general formula: ax^2 +bx + x in this case: a = 3 b = -12 sqrt(3) c = 36

OpenStudy (amistre64):

i typoed my formula :) ax^2 + bx + c

OpenStudy (amistre64):

we can simplfy these as well and still get the same answer by dividing off a 3 from them

OpenStudy (amistre64):

a = 3 b = -12 sqrt(3) c = 36 a = 1 b = -4 sqrt(3) c = 12

OpenStudy (anonymous):

do i go 1x^2+-4sqrt3+12

OpenStudy (amistre64):

not if you are trying to actuall use the quadratic formula itself.

OpenStudy (amistre64):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok then how do i do this

OpenStudy (amistre64):

plug in the values for a b and c into that formula and math away on it

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

just to give you something to go by for reference: \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\frac{-(-4\sqrt{3})\pm\sqrt{(-4\sqrt{3})^2-4(1)(12)}}{2(1)}\]

OpenStudy (anonymous):

whatdo i do after this point

OpenStudy (amistre64):

"Determine the exact values of the intercepts by using the quadratic formula. Then a calculator to evaluate the expressions that you obtain." That is where the math part come into the picture. all that adding, subtracting, multiplying and dividing amongst the numbers; and simplifying to get to a suitable answer .

OpenStudy (anonymous):

do i take the -(-4sqrt3)and do something with this at the begining?

OpenStudy (amistre64):

id make it positive and let it sit there till i did the messier part

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

hunh ... that messier part goes to zero and youre left with 4sqrt(3)/2; which reduce to 2sqrt(3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank for your patience